Whirly Corporation’s contribution format income statement for the most recent month is shown below:

Total Per Unit
Sales (8,100 units) $ 267,300 $ 33.00
Variable expenses 153,900 19.00
Contribution margin 113,400 $ 14.00
Fixed expenses 55,400
Net operating income $ 58,000

Required:

(Consider each case independently):

1. What would be the revised net operating income per month if the sales volume increases by 50 units?

2. What would be the revised net operating income per month if the sales volume decreases by 50 units?

3. What would be the revised net operating income per month if the sales volume is 7,100 units?
Answer #1
Case 1 Case 2 Case 3
Units sold (8100+50)=8150 (8100-50)=8050 7100
Sales (8150*33)=268950 (8050*33)=265,650 (7100*33)=234300
Variable expenses (8150*19)=154850 (8050*19)=152950 (7100*19)=134900
Contribution margin 114100 112700 99400
Fixed expenses 55400 55400 55400
Revised net operating income $58700 $57300 $44,000